### v2

# Consistency of Responses and Continuations Generated by Large Language Models on Social Media

Abstract
Large Language Models (LLMs) demonstrate remarkable capabilities in text generation, yet their emotional consistency and semantic coherence in social media contexts remain insufficiently understood. This study investigates how LLMs handle emotional content and maintain semantic relationships through continuation and response tasks using two open-source models: Gemma and Llama. By analyzing climate change discussions from Twitter and Reddit, we examine emotional transitions, intensity patterns, and semantic similarity between human-authored and LLM-generated content. Our findings reveal that while both models maintain high semantic coherence, they exhibit distinct emotional patterns: Gemma shows a tendency toward negative emotion amplification, particularly anger, while maintaining certain positive emotions like optimism. Llama demonstrates superior emotional preservation across a broader spectrum of affects. Both models systematically generate responses with attenuated emotional intensity compared to human-authored content and show a bias toward positive emotions in response tasks. Additionally, both models maintain strong semantic similarity with original texts, though performance varies between continuation and response tasks. These findings provide insights into LLMs’ emotional and semantic processing capabilities, with implications for their deployment in social media contexts and human-AI interaction design.
Introduction
Large Language Models (LLMs) represent one of the most significant yet controversial technological advancements in recent years. These models demonstrate unprecedented and expanding human-like capabilities, particularly in text generation, enabling diverse applications including text summarization (van Schaik and Pugh 2024), translation (Sung et al. 2024), and news writing (Muñoz-Ortiz, Gómez-Rodríguez, and Vilares 2024). Consequently, LLM-based applications have proliferated across domains, from conversational agents (Dam et al. 2024) to educational assistants (Liu, Jiang, and Wei 2025).
Despite their advantages, LLMs raise significant concerns regarding potential negative implications. These include content fabrication, commonly termed “hallucination,” which contributes to misinformation propagation (Huang et al. 2023). Furthermore, research indicates that LLM-generated content perpetuates societal biases encountered during training, potentially exacerbating AI fairness issues (Gallegos et al. 2024; Ayoub et al. 2024). Additionally, LLMs can influence human decision-making processes, potentially leading to unintended consequences through emotional manipulation or deception (Park et al. 2024). Given their widespread deployment, careful evaluation of LLMs’ text generation capabilities becomes imperative.
LLMs exhibit both task-specificity and context-sensitivity, with performance varying across different applications and contextual settings (Sung et al. 2024; Li, Zhang, and Sun 2023). Consequently, evaluating their text generation capabilities within realistic, socially relevant contexts becomes crucial. Social media platforms, serving as extensive networks for information exchange, provide valuable digital artifacts for such investigations.
In social media contexts, LLM text generation manifests in two primary forms: response tasks (e.g., replies) and continuation tasks (e.g., summarization and dialogue). The generated content influences public perception and engagement on social media platforms. Emotion embedded within text plays a crucial role as it can be rapidly activated and disseminated through extensive social networks, potentially facilitating emotional contagion (Kramer, Guillory, and Hancock 2014). Consequently, emotion serves as a strategic tool for engagement and persuasion in social media environments (Stieglitz and Dang-Xuan 2013; Hamby and Jones 2022).
Previous investigations of emotional effects on social media have employed real-life digital experiments through content manipulation (Kramer, Guillory, and Hancock 2014). However, such methodologies raise ethical concerns regarding unauthorized manipulation of user content and have generated public discomfort (Boyd 2016). LLMs offer a more ethically sound approach to examining emotional dynamics in social contexts. Given their increasingly sophisticated human-like capabilities, LLMs are extensively employed in simulating social interactions (Gao et al. 2024a). This LLM-based simulation methodology presents two key advantages: (1) AI agents can serve as safer substitutes for human participants in extreme or sensitive scenarios, and (2) they enable more controlled experimental conditions, facilitating precise examination of relevant variables.
With the widespread adoption of LLMs in generating human-like content, it becomes imperative to understand the consistency of LLM-generated text and its potential societal impact. Accordingly, this study investigates LLM text generation tasks (response generation and continuation) through systematic analysis of emotional consistency and semantic similarity. By examining these dynamics within climate change communication—a highly polarized and emotionally charged domain—this research addresses the following questions:
RQ1: How consistent are the emotions expressed in text generated by LLMs on social media?
RQ2: How does the emotional intensity of text generated by LLMs compare to text on social media?
RQ3: To what extent do LLMs demonstrate semantic similarity between generated text and text on social media?
By answering the research questions, this study has the following contributions:
  • Although LLMs demonstrate remarkable human-like capabilities in text generation, the mechanisms underlying their outputs remain insufficiently understood. Through analysis of emotional consistency between human- and LLM-generated text in social media contexts (response and continuation tasks), this study provides a deeper, contextualized understanding of LLMs’ text generation performance. The identified distinctions between human- and LLM-generated content illuminate how these models navigate social media interactions, which prompts critical ethical considerations regarding AI’s role in human interaction.
  • Emotion represents a fundamental behavioral response and crucial element in social media discourse. While previous research has predominantly focused on human-human communication, raising ethical concerns (Ferrara and Yang 2015), this study implements an LLM-based simulation approach. This methodology replicates human-AI agent interactions while addressing ethical limitations inherent in traditional research approaches, offering an innovative and ethically sound framework for emotion research.
  • Through analysis of real-world social media datasets concerning controversial scientific topics such as climate change, this study simulates scenarios where LLM-enabled tools participate in public discourse, presenting both opportunities and challenges (Feng et al. 2024). These findings enhance our understanding of communication dynamics and catalyze discussions regarding LLMs’ role in emotional guidance within controversial scientific discourse on social media platforms.
Related Works
Evaluation of LLMs generated text
The evaluation of LLM-generated text originates from natural language generation (NLG), defined as the process of computationally producing human-comprehensible text (Sai, Mohankumar, and Khapra 2022). Given the widespread deployment of AI models in text generation, extensive research has explored effective evaluation frameworks for NLG (Sai, Mohankumar, and Khapra 2022). Traditional evaluation metrics, primarily focused on quantifying content overlap between system outputs and references (Gao et al. 2024b), such as BLEU (Papineni et al. 2002) and ROUGE (Lin 2004), have served as standard metrics for automatically assessing output quality in machine translation and summarization tasks. However, these metrics demonstrate limitations when applied to complex, context-dependent tasks, particularly in the current generative AI paradigm (Gao et al. 2024b). Consequently, researchers have developed novel benchmarks for task-specific LLM evaluation (e.g., (Que et al. 2024)), while recent studies have proposed methodologies leveraging LLMs themselves for evaluation purposes (see (Gao et al. 2024b) for a comprehensive review).
The evaluation of LLM-generated text consistency with human behavior represents a fundamental approach to assessing model performance. Alignment with human behavior and response patterns remains a central objective in artificial intelligence development (Russell and Norvig 2016). Consistency is crucial for operational reliability and safety of LLMs, ensuring they can generate contextually appropriate and relatable outputs. Additionally, semantic similarity serves as an established metric for quantifying textual consistency (Chandrasekaran and Mago 2021). Researchers have evaluated LLM output consistency through semantic similarity measures and developed enhancement strategies to improve human alignment (Yang et al. 2024; Raj et al. 2023).
Existing literature predominantly examines distinctive characteristics between LLM- and human-generated text. For instance, (Herbold et al. 2023) conducted comparative analyses of human-written versus ChatGPT-generated essays across dimensions including topical coverage, logical structure, vocabulary usage, and linguistic constructions through human assessment. Beyond manual annotation, (Guo et al. 2023) implemented a mixed-methods approach to analyze LLM/human-generated responses across linguistic dimensions, revealing that LLM outputs demonstrate enhanced logical coherence, comprehensive detail, and reduced bias. (Muñoz-Ortiz, Gómez-Rodríguez, and Vilares 2024) employed quantitative analysis to compare human- and LLM-authored news content across morphological, syntactic, psychometric, and sociolinguistic dimensions. Through automated analysis, (Zanotto and Aroyehun 2024) identified distinctive linguistic patterns in text length, variability, syntactic complexity, and lexical diversity.
Text generation on social media context
In the social media environment, LLM text generation offers significant applications, including AI-powered social bots for online discourse participation, discussion summarization tools, and related applications (Li et al. 2024). However, ensuring generated text consistency requires careful consideration of contextual factors and interaction objectives. Social media interactions encompass both response generation (e.g., comment replies) and content continuation (e.g., social bot engagement). While existing research provides empirical evidence comparing human and LLM-generated content, the evaluation of social media-specific tasks, particularly responses and continuations, warrants comprehensive evaluation to understand LLM text generation in dynamic social media contexts.
Although emotion serves as a crucial factor in social media engagement and persuasion, its utilization as an evaluative feature for text generation remains insufficiently explored. Current comparative studies of human and LLM-generated text focus predominantly on static contexts, overlooking emotional dynamics. For instance, comparative analysis of human-written versus LLM-generated news content revealed stronger negative emotional expression in human-authored texts (Muñoz-Ortiz, Gómez-Rodríguez, and Vilares 2024). Similarly, while (Guo et al. 2023) examined response differences through multilingual sentiment classification, this approach presents limitations for comprehensive emotional analysis (e.g., distinguishing between joy and sadness). Given the dynamic nature of social media interactions, evaluation of emotional consistency in information exchange becomes crucial.
Emotional content permeates social media discourse and functions as a crucial determinant in shaping public opinion (Naskar et al. 2020). Emotion demonstrates high susceptibility to influence and serves as a critical factor in controversial and uncertain social agendas, including epidemics (Lu and Hong 2022), disasters (Chu et al. 2024), and polarizing social issues such as climate change (Brady et al. 2017). In these contexts, emotional responses can exert both beneficial and detrimental effects on public discourse. Climate change discourse, in particular, represents an extensively studied yet remains highly polarized domain, characterized by persistent denialism and skepticism (Treen, Williams, and O’Neill 2020; Whitmarsh 2011). These misconceptions frequently leverage emotional appeals, particularly fear, to influence public perception (Martel, Pennycook, and Rand 2020). Clinical psychology research has established correlations between anger, elevated cortisol responses to stress, and increased vulnerability to misinformation (Sharma, Wade, and Jobson 2023).
Above all, evaluating emotional patterns across response and continuation tasks within climate change discussions on social media provides a crucial framework for comparing LLM and human-generated content in dynamic, real-world scenarios.
Methodology
Experimental Design
Figure 1 illustrates the overall setup of the experimental design. In our experiment, we used two open-source large language models: Gemma [^1] https://ai.google.dev/gemma and Llama [^2] https://ai.meta.com/llama/license/, developed by Google and Meta, respectively. Specifically, we chose Gemma2-27B-Instruct-Q8 and Llama3-70B-Instruct, respectively, both of which excel in both performance and robustness. We utilized Ollama [^3] https://ollama.com/ as a framework to enable the two open source models to run on our local server.
Ollama supports various functions such as model creation, content generation, chat and embedding calculation. In our study, we mainly used the chat and embedding calculation. For the response task, we called its chat function directly, and unlike normal content generation functions, it allows the large language model to talk directly to the input content, which means there is no limitation of prompt words. This enables the most intuitive observation of the state of the large language model as an interlocutor, yielding more realistic and direct data. When it comes to the continuation task, we used the content generation function, which needs a prompt to ask the LLMs to expand the text it received. Again, here we made our prompts as concise as possible to minimize the impact of the prompts on the model. We tell the LLMs that “Assuming you are the author of this text, stand in your shoes and continue to expand the passage as you understand it.”
Figure 1: Experimental pipeline of consistency evaluation for LLMs. Our experimental framework begins with human text input to two LLMs (Gemma and Llama), which perform two distinct tasks: continuation and response. The continuation task employs a specific prompt instructing the model to expand the text as its author, while the response task operates without explicit prompting to enable natural interaction. Following content generation, we implement emotion detection on the outputs, followed by comprehensive analyses. The framework concludes with parallel analyses of emotional content and semantic similarity to evaluate the consistency of LLM-generated content relative to the original human input.
Dataset
This study utilized climate change corpora collected from Twitter (now X) and Reddit. We collected data using the Twitter Search API by querying relevant keywords, including climate change”, climate science”, climate manipulation”, climate Engineering”, climate Hacking”, climate modification”, Global Warming”, carbon footprint”, and “The Paris Agreement”. For Reddit, we used data maintained by Pushshift from https://the-eye.eu/redarcs/. The Pushshift Reddit dataset consists of two sets of files: submissions and comments(Baumgartner et al. 2020). The same keywords were applied to filter Reddit data, and to compare the differences in emotions, we collected both posts and comments from both platforms.
a
b
Figure 2: Daily data amount of Twitter and Reddit. a. Daily comments count of Reddit. b. Daily tweets count of Twitter. The x-axis represents the date, and the y-axis represents the frequency.
With the keywords, we obtained 5,768,822 Reddit comments and 76,596,654 tweets from Twitter. We used histograms to understand the basic distribution of data (Figure 2). To ensure temporal representation and minimize the impact of special events, we employed proportional sampling, selecting 200 rows per month systematically from Twitter data and 100 rows per month from Reddit data. This sampling strategy yielded a final dataset comprising 12,200 rows of Twitter data and 10,900 rows of Reddit data.
Emotion Labeling
In this study, we developed a methodology to analyze emotions in cross-platform social media data using a deep neural network-based model. We employed the twitter-roberta-base-emotion-multilabel-latest [^4] https://huggingface.co/cardiffnlp/twitter-roberta-base-emotion-multilabel-latest model from Hugging Face to examine the emotional content of both original texts and content generated by large language models. This model, built upon the RoBERTa-base architecture, is a fine-tuned version of “cardiffnlp/twitter-roberta-base-2022-154m” optimized on the SemEval 2018 Task 1 dataset—Affect in Tweets.
The twitter-roberta-base-emotion-multilabel-latest model identifies eleven distinct emotion categories: anticipation, joy, love, optimism, surprise, trust, anger, disgust, fear, pessimism, and sadness. The model outputs probability scores for each category, which serve as quantitative measures of emotional content for subsequent analysis. While previous studies used sentiment analysis (negative, positive, and neutral) for evaluating differences between human- and LLMs-generated text (Guo et al. 2023), our emotion-based approach provides a more granular and nuanced understanding of the underlying emotional states in posts.
Semantic Similarity
To assess both the comprehension capabilities of LLMs and their content generation accuracy, we employed cosine similarity as a quantitative measure of semantic alignment between LLMs’ outputs and human-generated texts. This methodology also enables the detection of semantic aberrations, commonly referred to as “hallucinations”, which frequently manifest in LLM outputs (Breum et al. 2023).
Cosine similarity, a fundamental metric in natural language processing, has demonstrated its utility across various text mining applications, including text classification, summarization, information retrieval, and question answering systems (Li and Han 2013). The mathematical foundation of this metric ensures robust comparison of textual similarities, making it particularly suitable for our analysis.
Due to the architectural differences between the two models, we implemented separate embedding calculations for LLM-generated and human-authored texts using both Gemma and Llama. The semantic similarity computations were performed using the cosine similarity function from the sklearn [^5] https://scikit-learn.org/stable/ library. For each platform, we derived four distinct similarity metrics quantifying the semantic relationships among four text pairs: Gemma’s continuation text, Gemma’s response text, Llama’s continuation text, and Llama’s response text, each compared against the original human text. To evaluate the comparative performance across different tasks and models, we employed the Mann-Whitney U-test to determine the statistical significance of variations in cosine similarity scores.
Results
Emotion Dynamics of the Original Text in Downstream Tasks
In this study, we examined the emotional transitions between human-generated text and LLM outputs in downstream tasks. We categorize 11 kind emotions into those that are positively oriented and those that are negatively oriented as followed(Robinson 2008) :
Positive emotions: anticipation, joy, love, optimism, surprise, trust(Vaillant 2008);
Negative emotions: anger, disgust, fear, pessimism, sadness
Figure 3: Emotional Transition Analysis of LLM Response and Continuation Tasks in Reddit Comments. Panels a, b, c, and d illustrate emotional transitions in content generated by Gemma and Llama models during continuation and response tasks, respectively. The y-axis represents source emotions from human text, while the x-axis indicates emotions in LLM-generated content. Cell values represent the proportion of emotional transitions between original and generated content. For example, in Figure 3a, the value 0.64 in the anger-to-anger cell indicates that 64% of originally angry texts maintained their emotional valence in Gemma’s continuation task. The intensity of each cell’s shading represents the proportion of emotional transition, with darker shades indicating higher transition frequencies.
Analysis of Figure 3a reveals that 62% of texts initially labeled as angry maintained their emotional valence in Gemma’s continuations. In contrast, only 11% of originally optimistic texts maintained their optimistic valence in the continuations. Other emotional categories, including anticipation, disgust, fear, joy, and sadness, predominantly shifted toward anger in the continuations, with transition rates of 29%, 44%, 30%, 39%, and 31%, respectively. Notably, optimism and surprise exhibited distinct patterns: 43% of optimistic texts preserved their emotional valence in the Gemma task, while texts expressing surprise demonstrated consistent emotional preservation. These findings suggest that the Gemma model exhibits a systematic tendency to transform diverse emotional expressions into anger during continuation tasks, indicating a bias toward negative emotional content, particularly anger. However, its performance with optimism and surprise demonstrates capacity for emotional preservation, suggesting selective ability to maintain certain emotional contexts throughout the generation process. In the analysis of emotional transitions during Gemma’s response task, we observed a significant shift in emotional valence, with more than 50% of initially angry texts transitioning toward anticipation and optimism in the responses.
Our analysis reveals that over 50% of texts with negative emotional valence transitioned toward positive expressions, particularly anticipation and optimism, in the response tasks. Texts with initial positive valence consistently preserved their affective characteristics, manifesting as anticipation, joy, or optimism. This pattern highlights Gemma’s systematic bias toward positive affect during response tasks. Significantly, a subset of original emotions still transitioned to anger in the responses, suggesting Gemma’s persistent sensitivity to anger-related content across both response and continuation tasks.
Examination of Llama 3’s performance, presented in Figure 3c and Figure 3d, demonstrates enhanced capability in emotional recognition and preservation during continuation tasks, maintaining original emotional valence with greater consistency.
The underlying mechanisms of emotional preservation in LLM-generated text appear to involve multiple factors: emotion-specific lexical choices, syntactic structures, and contextual integration during the generation process, coupled with the model’s capacity for sustaining emotional patterns. Emotions characterized by prominent semantic features (e.g., intense affective vocabulary) appear more readily preserved by the model.
Quantitative analysis of emotional transitions reveals that 21% of texts originally expressing anticipation, 48% expressing joy, and 17% expressing optimism transitioned to anger expressions. This pattern can be attributed to two primary factors: first, the semantic prominence of anger, characterized by explicit and intense features (e.g., critical, adversarial language) that are more readily recognized by the model during generation tasks; second, the contextual fragility of positive emotions, which typically require more complete contextual support, whereas LLM-generated content appears to favor more salient negative emotional expressions.
Consistent with Gemma’s response patterns, Llama’s response task demonstrated that, excluding fear, transitions to anticipation and optimism exceeded 50% across emotional categories. This consistent positive bias in response tasks across both models suggests a systematic emotional bias embedded during training, particularly evident in interactive contexts.
Figure 4: Emotional Transition Analysis of LLM Response and Continuation Tasks in Twitter Comments. Panels a, b, c, and d illustrate emotional transitions in content generated by Gemma and Llama models during continuation and response tasks on Twitter, respectively. The y-axis represents the original emotions in human-authored tweets, while the x-axis shows the emotions detected in LLM-generated content. Each cell value represents the proportion of emotional transitions, with darker shades of red indicating higher transition frequencies. For example, cell values of $1.0$ would indicate complete preservation of the original emotion, while lower values indicate transitions to different emotional states. The heatmap reveals patterns in how each model preserves or transforms emotional content across response and continuation tasks, with notable variations in the handling of positive (e.g., joy, optimism) versus negative (e.g., anger, fear) emotions.
In addition to Reddit, we analyzed data from Twitter, which exhibits distinct discourse patterns surrounding climate change. Figure 4 illustrates the emotional transitions between original Twitter content and LLM-generated responses. In Gemma’s continuation and response tasks, we observed that regardless of the original emotional valence—whether positive (anticipation, joy) or negative (anger, disgust)—the generated content predominantly expressed anger and anticipation. Among these transitions, anger represented the highest proportion, while anxiety emerged as the predominant emotion, followed by anticipation. A notable distinction between the two tasks was the increased frequency of responses transitioning toward anticipation. In Llama’s continuation task, the original emotional content similarly demonstrated a predominant shift toward anger. In contrast, Llama’s response task exhibited a primary shift toward anticipation. The performance patterns of both Gemma and Llama models across continuation and response tasks reveal two key insights: first, Gemma’s heightened sensitivity to anger-related content, and second, the models’ systematic bias toward positive affect when functioning in interactive dialogue contexts.
Resources of LLMs’ Generated Content Emotions
Figure 5: Emotional source analysis of LLM-generated content across platforms. Panels a, c, e, and g illustrate emotional transitions in Gemma and Llama models’ continuation and response tasks on Reddit data. Panels b, d, f, and h show corresponding results from Twitter data. Red lines indicate content generated from originally positive emotional text, while blue dashed lines represent content derived from negative emotional text. The absolute x-values represent the proportion of emotional source contributions in the generated content. The y-axis displays the emotional categories present in both original and generated content, while the x-axis ($-1.0$ to $1.0$) indicates the strength and direction of emotional transitions.
We analyze the emotional sources of LLM-generated content by examining the relationship between input and output emotions. In Gemma’s continuation task, Figure 5a reveals that positive emotions in generated content primarily derive from positive emotional sources. For instance, joy-labeled content originates from anticipation (32.04%), joy (25.24%), and optimism (13.59%). Conversely, content expressing negative emotions predominantly stems from negative emotional sources, with anger-labeled content derived primarily from original anger (61.8%) and disgust (13.03%). This pattern suggests Gemma’s tendency to maintain emotional valence consistency during content continuation.
Figure 5c illustrates Llama’s continuation task results, which differ from Gemma’s pattern. Most emotional content, except for optimism and joy, originates predominantly from negative emotional sources. While the preservation of negative emotional sources aligns with Gemma’s behavior, Llama distinctively generates positive emotional content primarily from negative emotional sources, potentially indicating an inherent positive affect bias.
Figure 5b,d illustrate the response tasks for Gemma and Llama, respectively. In Gemma’s responses, most positive emotional content, except joy, originates from predominantly negative emotional sources. Negative emotional content maintains its source valence, with anger-labeled content derived 69.55% from original anger expressions. Similarly, Llama’s responses demonstrate consistent transformation of negative emotions into positive ones, exemplified by anticipation-labeled content originating 37.23% from anger. However, negative emotional content maintains its original valence.
These findings demonstrate LLMs’ systematic transformation of negative emotions into positive ones during response tasks, while simultaneously exhibiting some degree of negative emotional preservation in their responses.
Comparative Analysis of Emotional Intensity between LLMs and Human Text
Beyond examining emotional transitions, we investigated the quantitative differences in emotional intensity between LLM-generated and human-authored content. This analysis specifically focused on determining whether LLM-generated content exhibits higher or lower emotional intensity compared to human expressions. To quantify emotional content, we employed a probabilistic model that assigns normalized scores (0 to 1) to each emotional category within the text. These probability values were interpreted as emotional intensity scores (Miyazaki et al. 2024). We categorized emotional expressions into five distinct intensity groups based on these scores. Statistical analysis comprised an ANOVA test to evaluate differences in emotional intensity across groups, followed by Tukey’s post-hoc test to identify significant pairwise variations(Elnaggar, Mohamed, and Gehan 2024).
Table 1: Test for significant differences in affective values between groups
Platform Emotions F statistic P value
Reddit anger 197.316 6.55E-166
anticipation 80.899 1.18E-67
disgust 35.307 4.12E-29
optimism 59.526 6.83E-50
fear 45.168 2.70E-37
sadness 18.736 3.58E-15
Twitter anger 384.285 9.5078E-320
anticipation 16092.479 0
disgust 179.181 1.80E-143
optimism 6.088 6.95E-05
fear 73.936 6.99E-60
joy 1742.163 0.00E+00
sadness 148.411 6.85E-98
Analysis of variance (ANOVA) results presented in Table 11 indicate statistically significant differences ($P<0.05$) across the five groups in emotional intensity values for anger, anticipation, disgust, optimism, fear, joy, and sadness on Twitter. Similar significant variations were observed in the Reddit dataset for anger, anticipation, disgust, optimism, fear, and sadness. Tukey’s post-hoc analysis identified several significant differences across three comparison categories: within-model, between-model, and model-to-human comparisons.
On Twitter discussions of climate change, within-model analyses revealed that Gemma’s continuation content exhibited significantly lower anticipation values compared to its response content. Similarly, Llama’s continuation content showed significantly lower anticipation values than its response content.
Between-model comparisons demonstrated that Gemma’s continuation content expressed significantly higher anticipation values than Llama’s continuation content. Additionally, Gemma’s response content showed significantly higher anticipation values compared to Llama’s responses. In model-to-human comparisons, both models’ generated content (continuation and response) demonstrated significantly lower anticipation values compared to the original human text.
These findings demonstrate that: (1) both models express higher anticipation values in response tasks compared to continuation tasks; (2) Gemma consistently generates content with higher anticipation values compared to Llama across both tasks; and (3) both models generate content with significantly reduced anticipation values compared to human-authored text, indicating a systematic reduction in emotional intensity during the generation process.
Table 2: Tukey’s post-hoc test of Reddit emotions
Emotion Within Group1 Between Groups2 Comparison with original
Gemma Con
vs. Resp
Llama Con
vs. Resp
Gemma Con
vs. Llama Con
Gemma Resp
vs. Llama Resp
Gemma Con
vs Original
Gemma Resp
vs Original
Llama Con
vs. Original
Llama Resp
vs. Original
Anticipate <*** <*** <*** <***
Joy <**
Disgust <*** <** >*** <*** >* <***
Sadness <*** >* <*** <*** <***
Anger <*** <*** <*** <*** <*** >***
Optimism >*** >** <*** <*** <*** <***
Fear <*** <*** <*** <*** <***
  • 1
    The comparison of the emotion values of continuation and response content generated by the same model.
  • 2
    The comparison of the emotion values of the same task but generated by different model.
  • 3
    The stars indicate the p values of the Mann–Whitney U test: *** for $p<0.001$, ** for $p<0.01$, and * for $p<0.05$.
  • 4
    The symbolic $<$ and $>$ indicate that the data in the previous column is less than or greater than the data in the next column. The “-” means there is no statistically significant difference between the two columns of data.
Table 3: Tukey’s post-hoc test of Twitter emotions
Emotion Within Group1 Between Groups1 Comparison with original1
Gemma Con
vs. Resp
Llama Con
vs. Resp
Gemma Con
vs. Llama Con
Gemma Resp
vs. Llama Resp
Gemma Con
vs Original
Gemma Resp
vs Original
Llama Con
vs. Original
Llama Resp
vs. Original
Anticipate <*** <*** >*** <*** <*** <*** <*** <***
Joy <*** <*** <*** <***
Disgust <*** <*** <* <*** <***
Sadness <*** <*** <*** <*** <*** <***
Anger <*** <*** <*** <*** <*** <*** <*** <***
Optimism <*** <*** <*** <***
Fear <*** <*** <** <** <*** <*** <*** <***
  • 1
    The comparison of the emotion values of continuation and response content generated by the same model.
  • 2
    The comparison of the emotion values of the same task but generated by different model.
  • 3
    The stars indicate the p values of the Mann–Whitney U test: *** for $p<0.001$, ** for $p<0.01$, and * for $p<0.05$.
  • 4
    The symbolic “$<$ and $>$” indicate that the data in the previous list is less than or greater than the data in the next list. The “-” means there is no statistically significant difference between the two columns of data.
Further analysis of additional emotional dimensions revealed that Tukey’s post hoc test results indicate significant differences ($P<0.05$) between LLM-generated and original texts across multiple emotional categories. Both Gemma and Llama’s generated content, in both continuation and response tasks, exhibited significantly lower intensities of positive emotions, particularly joy and optimism. A nuanced pattern emerged in the expression of negative emotions. The continuation texts generated by both models demonstrated attenuated levels of sadness, anger, disgust, and fear compared to their respective response texts and the original human content. Notably, Gemma’s response text exhibited significantly lower intensities of anger, disgust, and fear compared to Llama’s responses. These findings suggest two key insights: first, LLMs demonstrate systematic suppression of certain negative emotions, particularly in continuation tasks; second, the response task appears to operate under distinct generative mechanisms, resulting in differential emotional expression patterns. Furthermore, the consistent reduction in optimism across all LLM-generated texts relative to human-authored content indicates a systematic constraint in LLMs’ capability to fully capture and convey positive emotional states.
Semantic Coherence Analysis between LLM-Generated and Human-Authored Content
In LLM-human interactions, beyond examining emotional congruence and intensity patterns demonstrated in our previous experiments, we investigated the models’ capacity to maintain topical coherence and generate contextually relevant responses. To quantify this relationship, we employed cosine similarity as a metric to assess semantic alignment between generated and original content. This approach provides a quantitative framework for evaluating the semantic fidelity of LLM-generated responses across different interaction contexts.
a
b
Figure 6: Distribution of cosine similarity scores for LLM-generated content, visualized using violin plots comparing Gemma and Llama models’ continuations and responses across Reddit (panel a) and Twitter (panel b) platforms. Cosine similarity values range from -0.2 to 1.0, with median scores indicated numerically for each distribution.
Table 4: Mann-Whitney U test for cosine similarity between LLM-generated content and two platform text
Comparison
group
Statistics P value
Reddit Inner
Gemma
continuation
vs
Gemma
response
56740052 0.999999995
Llama
continuation
vs
Llama
response
66851679 4.05E-58
Inter
Gemma
continuation vs
Llama
continuation
68723025 8.90041E-90
Gemma
response vs
Llama
response
78292192 0.00E+00
Twitter Inner
Gemma
continuation vs
Gemma
response
93014920 9.74E-251
Llama
continuation vs
Llama
response
60081284 1
Inter
Gemma
continuation vs
Llama
continuation
118641175 0
Gemma
response vs
Llama
response
69518835 1
  • 1
    The $p<0.05$ indicate that the data in the previous group is significantly higher than the data in the latter group.
Table 4 presents the Mann-Whitney U test results across the four experimental conditions. The analysis reveals that both Gemma and Llama maintained substantial semantic alignment with the original text in their continuation and response tasks. Figure 6a illustrates that semantic similarity values on the Twitter platform predominantly exceeded 0.5, indicating strong semantic coherence between LLM-generated content and input text. This suggests the models’ capability to comprehend and generate contextually relevant responses while maintaining topical coherence. Similar patterns emerged in the Reddit dataset, as shown in Figure 6b, where all four sets of cosine similarities demonstrated values above 0.5.
The Mann-Whitney U test (Table 4) revealed significant within-model and cross-task variations in semantic similarity. On the Twitter platform, Gemma’s continuation task demonstrated significantly higher semantic similarity compared to its response task, while Llama exhibited the opposite pattern, with responses showing significantly higher similarity ($P<0.01$). Cross-model analysis revealed task-specific differences: Gemma’s continuation task demonstrated significantly higher semantic similarity compared to Llama’s continuation task ($P<0.01$), while Llama achieved higher similarity scores in the response task. On Reddit, Llama’s continuation content exhibited significantly higher semantic similarity compared to its response content ($P<0.01$). Furthermore, Gemma’s generated content demonstrated significantly higher semantic similarity compared to Llama’s in both continuation and response tasks ($P<0.01$).
Discussion
This study investigated the consistency between LLM-generated content and human input, with particular focus on emotional content transmission. Given that text often encodes substantial emotional information, we hypothesized that LLMs may demonstrate diverse emotional patterns. As LLMs increasingly integrate into human social contexts through chatbots and social agents, understanding their semantic consistency and emotional transmission patterns becomes crucial for human-machine interaction. We decomposed the fundamental question of LLM emotional capability into three specific research questions, based on our empirical findings. Several key insights emerged from our analysis.
In the continuation task, Gemma demonstrated a systematic transformation of most original emotions toward anger, indicating a bias toward negative emotional expression, particularly in anger intensification. However, Gemma exhibited capacity for emotional preservation, specifically for optimism and surprise, suggesting ability to perceive and maintain certain emotional valences through the continuation process. Conversely, the Llama model demonstrated enhanced capability in preserving original emotional states during continuation tasks, particularly for anger, anticipation, fear, optimism, and sadness, with relatively lower emotional transformation rates, indicating superior emotional continuity (RQ1).
We posit that during continuation tasks, LLMs employ emotion-specific lexical markers, syntactic patterns, and contextual cues to maintain emotional consistency with source texts. Emotions characterized by prominent semantic features (such as anger, anticipation, and sadness) typically manifest through distinct emotional vocabulary and strong affective markers, facilitating model recognition and continuation. Moreover, the semantic salience of anger may enhance its recognition and preservation in generative tasks, while positive emotions like optimism or joy may require more robust contextual support, making them susceptible to disruption by LLM-generated negative content.
In the response task, both Gemma and Llama models exhibited a systematic bias toward positive emotions, predominantly converting original emotions into anticipation and optimism. This pattern suggests an embedded bias from the training process favoring positive affect. However, the persistent presence of anger in a subset of responses demonstrates these models’ susceptibility to negative emotional content during the response generation process.
Our experimental findings indicate that both Gemma and Llama models exhibit emotional capabilities and can effectively recognize and maintain human emotional states in continuation tasks. This emotional recognition capability is fundamental for LLMs, as it underlies their ability to comprehend both user requirements and emotional context.
Regarding emotional intensity (RQ2), our results demonstrate that LLM-generated content exhibits significantly attenuated intensity across most emotional dimensions compared to human-authored texts. This suggests that LLMs exhibit more moderated emotional expression, maintaining more constrained emotional ranges across both positive and negative affects. Furthermore, the consistently lower emotional intensity values in both continuation and response tasks, compared to original texts, indicates potential limitations in LLMs’ capacity to fully capture and convey emotional depth.
Analysis of emotional types reveals distinct patterns in LLM-generated content. Regarding positive emotions, LLM-generated texts demonstrate significantly lower intensities of joy and optimism compared to human-authored texts, suggesting limitations in the models’ capacity to fully comprehend and express positive emotional states characteristic of human experience. Similarly, for negative emotions, LLM-generated content exhibits attenuated intensities of sadness, anger, disgust, and fear, indicating systematic suppression of negative emotional expression in downstream tasks.
In the context of global challenges such as climate change, which transcends temporal, spatial, and geographical boundaries and impacts long-term human development, public discourse exhibits diverse emotional responses. As social media platforms rapidly evolve, they become crucial channels for public communication during climate-related events, such as the Australian bushfires. The effective management of public emotions during critical periods following such events represents a key component of public opinion governance. Recent advances in Artificial Intelligence Generated Content (AIGC), driven by Generative AI (GAI) technology, have garnered attention beyond computer science (Cao et al. 2023). Given the increasing integration of LLMs into daily life, their emotional characteristics significantly influence opinion leadership, as emotional content shapes public perception and discourse framing.
Our experimental results (Figure 3 and Figure 4) demonstrate that LLMs systematically transition toward positive emotions during response tasks. This positive bias suggests potential utility in public opinion events, offering constructive responses and mitigating negative public sentiment. However, the observed patterns of negative emotional transmission pose potential societal risks. LLMs’ capability to generate contextually relevant content could be exploited for malicious purposes. Beyond concerns regarding misinformation dissemination, the potential intensification of emotional polarization presents a significant challenge in LLM deployment.
From the perspective of human-computer interaction, emotional support serves as a fundamental component in enhancing social interactions, facilitating psychological interventions, and improving customer service outcomes through addressing emotional needs. The quality of emotional support and user understanding significantly impacts long-term user engagement and trust in LLM interactions (Schneider, Flores, and Kranz 2024).
Addressing our third research question, we analyzed embedding representations of generated and original content, employing cosine similarity as a metric for semantic proximity. Our findings indicate that both Gemma and Llama consistently demonstrate high cosine similarity values, suggesting robust capability in capturing and reproducing semantic features. Models have shown a remarkable ability to represent, comprehend, and generate human-like text. Compared to prior Natural Language Processing (NLP) approaches, one of the most striking advances of LLMs is their ability to generalize their “knowledge” to novel scenarios, contexts, and tasks(Peters and Matz 2024) Future implications suggest that while LLMs demonstrate competence in generating semantically coherent content, opportunities exist for enhancing alignment between generated outputs and nuanced human context. Future research directions should explore mechanisms for refining LLMs’ comprehension of implicit meaning and contextual subtleties, thereby enhancing user experience and expanding application domains.
Limitation and Future Work
This study contributes to understanding the emotional dynamics of human-AI interactions, while suggesting several avenues for future research based on current limitations.
First, regarding experimental design, our study was limited to data from Reddit and Twitter platforms, potentially under-representing the broader social media ecosystem. Other platforms such as YouTube, Instagram, and TikTok exhibit distinct user behaviors and content structures (Hilde A. M. Voorveld and Bronner 2018). Furthermore, our analysis was bounded by the capabilities of open-source emotion models (Llama and Gemma), which may demonstrate different performance characteristics compared to proprietary commercial models like ChatGPT or Google’s Bard. Future research could expand the scope by incorporating data from additional social media platforms and evaluating commercial models to enhance generalizability.
Second, our experimental scope, focused on English-language climate change discourse, provides limited insight into model performance across diverse topics and languages. The emotional expression patterns of LLMs may vary significantly across different subject domains. Future research directions could incorporate multilingual datasets to identify cross-linguistic semantic variations (Zhao et al. 2020) and investigate potential biases within LLMs across isolated or intersecting semantic spaces (Hassan et al. 2018).
Third, while we utilized social media to obtain human-authored content, the dataset potentially includes automatically generated content and social bot interactions that may not accurately represent genuine human expression. Future research should implement more rigorous data validation protocols to enhance dataset quality and control for confounding variables.
Conclusion
This study examined the semantic and emotional consistency of content generated by LLMs. Recognizing the emotional information embedded in text and the way LLMs handle these emotions is essential as they become more involved in human social contexts.
Our findings revealed that Gemma demonstrated a tendency to amplify negative emotions, particularly anger. Despite this bias, Gemma effectively perceived and transferred the original emotional tone of the text. On the other hand, the Llama model displayed stronger emotional retention across a broader spectrum, including anger, expectation, fear, optimism, and sadness, with fewer transitions to other emotions. For the semantic problem, both models show better performance for both continuation and response. LLMs can understand human input to some extent. These observations highlight distinct emotional dynamics between LLMs, offering insights for enhancing their design and application in emotion-sensitive contexts.
Acknowledgments
This study is partially supported by the 2023-2024 Annual Seed Research Funding of the Department of Science and Technology of Communication, University of Science and Technology of China (#ASRF2023).

### v3

@@
 Large Language Models (LLMs) demonstrate remarkable capabilities in text generation, yet their emotional consistency and semantic coherence in social media contexts remain insufficiently understood.
-This study investigates how LLMs handle emotional content and maintain semantic relationships through continuation and response tasks using two open-source models: Gemma and Llama.
-By analyzing climate change discussions from Twitter and Reddit, we examine emotional transitions, intensity patterns, and semantic similarity between human-authored and LLM-generated content.
-Our findings reveal that while both models maintain high semantic coherence, they exhibit distinct emotional patterns: Gemma shows a tendency toward negative emotion amplification, particularly anger, while maintaining certain positive emotions like optimism.
-Llama demonstrates superior emotional preservation across a broader spectrum of affects.
-Both models systematically generate responses with attenuated emotional intensity compared to human-authored content and show a bias toward positive emotions in response tasks.
-Additionally, both models maintain strong semantic similarity with original texts, though performance varies between continuation and response tasks.
-These findings provide insights into LLMs’ emotional and semantic processing capabilities, with implications for their deployment in social media contexts and human-AI interaction design.
+This study investigates how LLMs handle emotional content and maintain semantic relationships through continuation and response tasks using three open-source models: Gemma, Llama3 and Llama3.3 and one commercial Model:Claude.
+By analyzing climate change discussions from Twitter and Reddit, we examine emotional transitions, intensity patterns, and semantic consistency between human-authored and LLM-generated content.
+Our findings reveal that while both models maintain high semantic coherence, they exhibit distinct emotional patterns: these models show a strong tendency to moderate negative emotions.
+When the input text carries negative emotions such as anger, disgust, fear, or sadness, LLM tends to generate content with more neutral emotions, or even convert them into positive emotions such as joy or surprise.
+At the same time, we compared the LLM-generated content with human-authored content.
+The four models systematically generated responses with reduced emotional intensity and showed a preference for neutral rational emotions in the response task.
+In addition, these models all maintained a high semantic similarity with the original text, although their performance in the continuation task and the response task was different.
+These findings provide deep insights into the emotion and semantic processing capabilities of LLM, which are of great significance for its deployment in social media environments and human-computer interaction design.
 Introduction
@@
 Consequently, emotion serves as a strategic tool for engagement and persuasion in social media environments (Stieglitz and Dang-Xuan 2013; Hamby and Jones 2022).
-Previous investigations of emotional effects on social media have employed real-life digital experiments through content manipulation (Kramer, Guillory, and Hancock 2014).
-However, such methodologies raise ethical concerns regarding unauthorized manipulation of user content and have generated public discomfort (Boyd 2016).
-LLMs offer a more ethically sound approach to examining emotional dynamics in social contexts.
-Given their increasingly sophisticated human-like capabilities, LLMs are extensively employed in simulating social interactions (Gao et al.
-2024a).
-This LLM-based simulation methodology presents two key advantages: (1) AI agents can serve as safer substitutes for human participants in extreme or sensitive scenarios, and (2) they enable more controlled experimental conditions, facilitating precise examination of relevant variables.
+Previous investigations of emotional effects on social media have employed real-life experiments through content manipulation (Kramer, Guillory, and Hancock 2014).
+However, it raise ethical concerns regarding manipulation of user content and public discomfort (Boyd 2016).
+Recent work demonstrates that LLM-based simulations can be a useful tool for modeling social interactions (Gao et al.
+2024a): (1) AI agents can serve as safer substitutes for human participants in extreme or sensitive scenarios, and (2) they enable more controlled experimental conditions, facilitating precise examination of relevant variables.
 With the widespread adoption of LLMs in generating human-like content, it becomes imperative to understand the consistency of LLM-generated text and its potential societal impact.
+Moreover, with a growing variety of LLMs being adopted in real-world settings, there remains a lack of systematic, comparative evaluations of how consistent/different these models perform.
 Accordingly, this study investigates LLM text generation tasks (response generation and continuation) through systematic analysis of emotional consistency and semantic similarity.
@@
 RQ3: To what extent do LLMs demonstrate semantic similarity between generated text and text on social media?
-By answering the research questions, this study has the following contributions:
-  • Although LLMs demonstrate remarkable human-like capabilities in text generation, the mechanisms underlying their outputs remain insufficiently understood.
-Through analysis of emotional consistency between human- and LLM-generated text in social media contexts (response and continuation tasks), this study provides a deeper, contextualized understanding of LLMs’ text generation performance.
-The identified distinctions between human- and LLM-generated content illuminate how these models navigate social media interactions, which prompts critical ethical considerations regarding AI’s role in human interaction.
-  • Emotion represents a fundamental behavioral response and crucial element in social media discourse.
-While previous research has predominantly focused on human-human communication, raising ethical concerns (Ferrara and Yang 2015), this study implements an LLM-based simulation approach.
-This methodology replicates human-AI agent interactions while addressing ethical limitations inherent in traditional research approaches, offering an innovative and ethically sound framework for emotion research.
-  • Through analysis of real-world social media datasets concerning controversial scientific topics such as climate change, this study simulates scenarios where LLM-enabled tools participate in public discourse, presenting both opportunities and challenges (Feng et al.
-2024).
-These findings enhance our understanding of communication dynamics and catalyze discussions regarding LLMs’ role in emotional guidance within controversial scientific discourse on social media platforms.
+By answering the research questions, this study has the following findings and contributions:
+  • We systematically evaluated emotional and semantic consistency between LLM-generated and human-authored social media texts, finding that (1) LLMs’ emotional expressions differ significantly from original texts across both continuation and response tasks, and across different model types; (2) LLMs maintain high semantic similarity with original content, indicating strong capabilities in understanding and generating human-like text.
+  • We found that all models exhibit significantly lower emotional intensity in both tasks, suggesting that LLMs may struggle to convey the full emotional depth of human-authored content.
+  • We discussed the real-world implications of these emotional differences and semantic proximity — including how LLMs may influence emotional engagement, serve as tools for moderating polarized debates, or, conversely, be misused for emotional manipulation.
+  • We proposed an ethically grounded simulation framework using LLM agents to explore emotional dynamics around controversial topics on social media, avoiding privacy and consent issues inherent in traditional research (Ferrara and Yang 2015).
+We also adopted the “LLM-as-judge” approach to assess semantic similarity, reducing reliance on time-consuming and costly human annotations.
 Related Works
@@
 Figure 1 illustrates the overall setup of the experimental design.
-In our experiment, we used two open-source large language models: Gemma [^1] https://ai.google.dev/gemma and Llama [^2] https://ai.meta.com/llama/license/, developed by Google and Meta, respectively.
-Specifically, we chose Gemma2-27B-Instruct-Q8 and Llama3-70B-Instruct, respectively, both of which excel in both performance and robustness.
+Our core experiment involves using LLMs to interact with human-generated text data, which consists of social media posts collected from real users.
+In our experiment, we employed three open-source large language models—Gemma [^1] https://ai.google.dev/gemma, Llama 3222https://ai.meta.com/llama/license/, and Llama 3.3—developed by Google and Meta, respectively, as well as one commercial model, Claude 3.5, developed by Anthropic [^3] https://www.anthropic.com/claude.
+Specifically, we utilized the following model variants: Gemma2-27B-Instruct-Q8, Llama3-70B-Instruct, Llama3.3-70B, and Claude-3-5-Haiku-20241022-X.
+These models were selected due to their strong performance and robustness.
 We utilized Ollama [^4] https://ollama.com/ as a framework to enable the two open source models to run on our local server.
-Ollama supports various functions such as model creation, content generation, chat and embedding calculation.
-In our study, we mainly used the chat and embedding calculation.
-For the response task, we called its chat function directly, and unlike normal content generation functions, it allows the large language model to talk directly to the input content, which means there is no limitation of prompt words.
-This enables the most intuitive observation of the state of the large language model as an interlocutor, yielding more realistic and direct data.
-When it comes to the continuation task, we used the content generation function, which needs a prompt to ask the LLMs to expand the text it received.
-Again, here we made our prompts as concise as possible to minimize the impact of the prompts on the model.
-We tell the LLMs that “Assuming you are the author of this text, stand in your shoes and continue to expand the passage as you understand it.”
+In our study, we mainly used the chat of Ollama.
+For the response task, we directly used the chat interface, which allows LLMs to interact with the input content without requiring explicitly crafted instruction prompts.
+This setup offers a more natural and direct observation of the LLM’s behavior as a conversational agent.
+However, we acknowledge the reviewer’s point and clarify that although we did not provide any additional instruction prompt, the original human post itself serves as the input prompt to the model.
+For the continuation task, we used the content generation function, where concise prompts were employed to minimize their influence on the model, allowing for a clearer focus on the model’s response.We tell the LLMs that “Assuming you are the author of this text, stand in your shoes and continue to expand the passage as you understand it.”
 Figure 1: Experimental pipeline of consistency evaluation for LLMs.
-Our experimental framework begins with human text input to two LLMs (Gemma and Llama), which perform two distinct tasks: continuation and response.
+Our experimental framework begins with human text input to four LLMs , which perform two distinct tasks: continuation and response.
 The continuation task employs a specific prompt instructing the model to expand the text as its author, while the response task operates without explicit prompting to enable natural interaction.
 Following content generation, we implement emotion detection on the outputs, followed by comprehensive analyses.
-The framework concludes with parallel analyses of emotional content and semantic similarity to evaluate the consistency of LLM-generated content relative to the original human input.
+The framework concludes with parallel analyses of emotional content and semantic consistency to evaluate the consistency of LLM-generated content relative to the original human input.
 Dataset
 This study utilized climate change corpora collected from Twitter (now X) and Reddit.
-We collected data using the Twitter Search API by querying relevant keywords, including climate change”, climate science”, climate manipulation”, climate Engineering”, climate Hacking”, climate modification”, Global Warming”, carbon footprint”, and “The Paris Agreement”.
+We collected data using the Twitter Search API by querying relevant keywords, including “climate change”, “climate science”, “climate manipulation”, “climate Engineering”, “climate Hacking”, “climate modification”, “Global Warming”, “carbon footprint”, and “The Paris Agreement”.
 For Reddit, we used data maintained by Pushshift from https://the-eye.eu/redarcs/.
@@
 We used histograms to understand the basic distribution of data (Figure 2).
-To ensure temporal representation and minimize the impact of special events, we employed proportional sampling, selecting 200 rows per month systematically from Twitter data and 100 rows per month from Reddit data.
-This sampling strategy yielded a final dataset comprising 12,200 rows of Twitter data and 10,900 rows of Reddit data.
+Twitter and Reddit are two major social media platforms that differ from each other.
+Twitter is designed to be open, concise, and immediate through short posts and real-time updates, while Reddit is a community-based medium designed for deep, deliberative, and topical discussions(Parsa et al.
+2022; Treen et al.
+2022a).
+These differences in platform design not only shape how content is generated and shared, but also attract different user groups with different interests and engagement styles, which provides a valuable resource for understanding different modes of public discussion(Ruan et al.
+2022).
+We support these distinctions with some quantitative evidence using the average word length, word cloud.
+The results show that the average length of Reddit text is 656.66 words, while the average length of Twitter text is 246.34 words.
+This discrepancy reflects fundamental differences in how the two platforms are used.
+Reddit is structured around topic-based discussions within subreddits, encouraging longer, more detailed posts that often resemble mini-essays or narratives.
+The word cloud chart (see the appendixLABEL:wordcloud) shows that the expressions of users on the two platforms under the same topic are also different.
+For example, more people on Reddit are discussing ”climate change”, while more people on Twitter use ”global warming”.
+To construct datasets representative of the discussion dynamics surrounding climate change over an extended period and to mitigate potential biases introduced by sudden climate change events, we employed a time-stratified sampling approach on the raw data.
+Specifically, based on the collected corpus of 5,768,822 Reddit comments and 76,596,654 Twitter tweets, we initiated the process by partitioning the data by month.
+Subsequently, within each monthly segment, we conducted systematic sampling, extracting 200 records from the Twitter data and 100 records from the Reddit data.
+This method of monthly proportional sampling was implemented to ensure a uniform distribution of the dataset across the temporal dimension, thereby reducing the influence of transient extreme events or fluctuations in topic popularity during specific timeframes.
+This strategy aims to enable the final dataset to more comprehensively reflect the content and evolving trends of discussions pertaining to climate change across different temporal stages.
+Ultimately, this process yielded an analytical dataset comprising 12,200 Twitter data points and 10,900 Reddit data points.
 Emotion Labeling
 In this study, we developed a methodology to analyze emotions in cross-platform social media data using a deep neural network-based model.
-We employed the twitter-roberta-base-emotion-multilabel-latest [^4] https://huggingface.co/cardiffnlp/twitter-roberta-base-emotion-multilabel-latest model from Hugging Face to examine the emotional content of both original texts and content generated by large language models.
-This model, built upon the RoBERTa-base architecture, is a fine-tuned version of “cardiffnlp/twitter-roberta-base-2022-154m” optimized on the SemEval 2018 Task 1 dataset—Affect in Tweets.
-The twitter-roberta-base-emotion-multilabel-latest model identifies eleven distinct emotion categories: anticipation, joy, love, optimism, surprise, trust, anger, disgust, fear, pessimism, and sadness.
+We employed thej-hartmann/emotion-english-distilroberta-base [^5] https://huggingface.co/j-hartmann/emotion-english-distilroberta-base model from Hugging Face to examine the emotional content of both original texts and content generated by large language models.
+This model, built upon the RoBERTa-base architecture, is a fine-tuned checkpoint of “DistilRoBERTa-base” the datasets contain emotion labels for texts from Twitter, Reddit, student self-reports, and utterances from TV dialogues.
+The emotion-english-distilroberta-base model identifies seven distinct emotion categories:joy, surprise, neutral, anger, disgust, fear and sadness.
 The model outputs probability scores for each category, which serve as quantitative measures of emotional content for subsequent analysis.
-While previous studies used sentiment analysis (negative, positive, and neutral) for evaluating differences between human- and LLMs-generated text (Guo et al.
+While previous studies used sentiment analysis (negative, positive, and neutral) for evaluating differences between human and LLMs-generated text (Guo et al.
 2023), our emotion-based approach provides a more granular and nuanced understanding of the underlying emotional states in posts.
-Semantic Similarity
-To assess both the comprehension capabilities of LLMs and their content generation accuracy, we employed cosine similarity as a quantitative measure of semantic alignment between LLMs’ outputs and human-generated texts.
-This methodology also enables the detection of semantic aberrations, commonly referred to as “hallucinations”, which frequently manifest in LLM outputs (Breum et al.
+Semantic Consistency
+We design a evaluation framework that uses a rubric-guided expert-like LLM to score consistency and faithfulness.
+Here we hired the model:gemini-2.0-flash-thinking-exp-1219 as our expert LLM.
+This allows us to directly assess whether a model-generated response is logically aligned with the user’s comment, and please see the appendix for specific prompt designLABEL:lst:build-prompt.
+After obtaining the evaluation results of the LLM-as-judge, the results are reviewed by a human evaluator.
+Human evaluator judges whether the evaluation results of the LLM are reasonable and accurate.
+In the end, the consistency rate between the two is maintained at 75%.
+Automatic evaluations of instruction following abilities in LLMs has recently received significant attention (Zheng et al.
 2023).
-Cosine similarity, a fundamental metric in natural language processing, has demonstrated its utility across various text mining applications, including text classification, summarization, information retrieval, and question answering systems (Li and Han 2013).
-The mathematical foundation of this metric ensures robust comparison of textual similarities, making it particularly suitable for our analysis.
-Due to the architectural differences between the two models, we implemented separate embedding calculations for LLM-generated and human-authored texts using both Gemma and Llama.
-The semantic similarity computations were performed using the cosine similarity function from the sklearn [^5] https://scikit-learn.org/stable/ library.
-For each platform, we derived four distinct similarity metrics quantifying the semantic relationships among four text pairs: Gemma’s continuation text, Gemma’s response text, Llama’s continuation text, and Llama’s response text, each compared against the original human text.
-To evaluate the comparative performance across different tasks and models, we employed the Mann-Whitney U-test to determine the statistical significance of variations in cosine similarity scores.
+Given the significant time and cognitive effort required for human evaluation of large-scale generated content, we adopted a stratified random sampling approach to ensure representativeness across different model-platform-task combinations.
+Our full dataset consists of different sub-corpora, each corresponding to a specific combination of model (e.g., Gemma, Llama 3, Llama 3.3, Claude), platform (Twitter, Reddit), and task type (response or continuation), with approximately 10,000 instances per sub-corpus.
+To balance coverage and feasibility, we randomly sampled 30 instances from each sub-corpus, resulting in a total of 480 samples for evaluation.
+Manual expert evaluation is costly.
+This hybrid evaluation strategy combines manual judgment with the assistance of a LLM.
+Although the sample size is small, it covers diversity across platforms, tasks, and models.
+We acknowledge in the limitations section that this small sample size may affect the generalizability of our findings.
 Results
@@
 In this study, we examined the emotional transitions between human-generated text and LLM outputs in downstream tasks.
-We categorize 11 kind emotions into those that are positively oriented and those that are negatively oriented as followed(Robinson 2008) :
+We categorize 7 emotions as positive emotionally oriented and negatively oriented as well as neutral emotions as defined by itself, as follows (Robinson 2008) :
 Positive emotions: anticipation, joy, love, optimism, surprise, trust  (Vaillant 2008);
@@
 Figure 3: Emotional Transition Analysis of LLM Response and Continuation Tasks in Reddit Comments.
-Panels a, b, c, and d illustrate emotional transitions in content generated by Gemma and Llama models during continuation and response tasks, respectively.
+Panels a, b, c, d, e, f, g, and h illustrate emotional transitions in content generated by Gemma, Llama and Claude models during continuation and response tasks, respectively.
 The y-axis represents source emotions from human text, while the x-axis indicates emotions in LLM-generated content.
 Cell values represent the proportion of emotional transitions between original and generated content.
-For example, in Figure 3a, the value 0.64 in the anger-to-anger cell indicates that 64% of originally angry texts maintained their emotional valence in Gemma’s continuation task.
+For example, in Figure 3a, the value 0.34 in the anger-to-anger cell indicates that 34% of originally angry texts maintained their emotional valence in Gemma’s continuation task.
 The intensity of each cell’s shading represents the proportion of emotional transition, with darker shades indicating higher transition frequencies.
-Analysis of Figure 3a reveals that 62% of texts initially labeled as angry maintained their emotional valence in Gemma’s continuations.
-In contrast, only 11% of originally optimistic texts maintained their optimistic valence in the continuations.
-Other emotional categories, including anticipation, disgust, fear, joy, and sadness, predominantly shifted toward anger in the continuations, with transition rates of 29%, 44%, 30%, 39%, and 31%, respectively.
-Notably, optimism and surprise exhibited distinct patterns: 43% of optimistic texts preserved their emotional valence in the Gemma task, while texts expressing surprise demonstrated consistent emotional preservation.
-These findings suggest that the Gemma model exhibits a systematic tendency to transform diverse emotional expressions into anger during continuation tasks, indicating a bias toward negative emotional content, particularly anger.
-However, its performance with optimism and surprise demonstrates capacity for emotional preservation, suggesting selective ability to maintain certain emotional contexts throughout the generation process.
-In the analysis of emotional transitions during Gemma’s response task, we observed a significant shift in emotional valence, with more than 50% of initially angry texts transitioning toward anticipation and optimism in the responses.
-Our analysis reveals that over 50% of texts with negative emotional valence transitioned toward positive expressions, particularly anticipation and optimism, in the response tasks.
-Texts with initial positive valence consistently preserved their affective characteristics, manifesting as anticipation, joy, or optimism.
-This pattern highlights Gemma’s systematic bias toward positive affect during response tasks.
-Significantly, a subset of original emotions still transitioned to anger in the responses, suggesting Gemma’s persistent sensitivity to anger-related content across both response and continuation tasks.
-Examination of Llama 3’s performance, presented in Figure 3c and Figure 3d, demonstrates enhanced capability in emotional recognition and preservation during continuation tasks, maintaining original emotional valence with greater consistency.
-The underlying mechanisms of emotional preservation in LLM-generated text appear to involve multiple factors: emotion-specific lexical choices, syntactic structures, and contextual integration during the generation process, coupled with the model’s capacity for sustaining emotional patterns.
-Emotions characterized by prominent semantic features (e.g., intense affective vocabulary) appear more readily preserved by the model.
-Quantitative analysis of emotional transitions reveals that 21% of texts originally expressing anticipation, 48% expressing joy, and 17% expressing optimism transitioned to anger expressions.
-This pattern can be attributed to two primary factors: first, the semantic prominence of anger, characterized by explicit and intense features (e.g., critical, adversarial language) that are more readily recognized by the model during generation tasks; second, the contextual fragility of positive emotions, which typically require more complete contextual support, whereas LLM-generated content appears to favor more salient negative emotional expressions.
-Consistent with Gemma’s response patterns, Llama’s response task demonstrated that, excluding fear, transitions to anticipation and optimism exceeded 50% across emotional categories.
-This consistent positive bias in response tasks across both models suggests a systematic emotional bias embedded during training, particularly evident in interactive contexts.
+Analysis of Figure 3a reveals that in Gemma’s continuation tasks, 34% of the texts initially labeled as ’anger’ maintained their emotional, and 38% of the texts originally categorized as ’fear’ also preserved the ’fear’ emotion in the continuation task.
+For other emotion categories, including ’disgust,’ ’sadness,’ ’neutral,’ ’joy,’ and ’surprise,’ the proportions of texts that retained the same emotional label were 22%, 13%, 40%, 10%, and 1%, respectively, demonstrating a capability for emotion preservation.
+It is noteworthy that for all original emotion labels, a considerable proportion of the texts shifted to a neutral label in the continuation task.
+Concurrently, a significant portion of texts also transitioned to the ’anger’ label during the continuation.
+These findings suggest that the Gemma model possesses a certain ability to recognize and perpetuate the original emotions.
+In subsequent tasks, it exhibits a systematic tendency to convert various emotional expressions into a neutral sentiment.
+The results in Figure 3a also, to some extent, indicate the model’s sensitivity to negative emotions in the continuation task, particularly anger.
+In analyzing the emotional shifts within Gemma’s response tasks, we observed significant changes in emotional conversion.
+For texts with originally positive emotions such as ’joy’ and ’surprise,’ as well as those with negative emotions like ’anger,’ ’disgust,’ ’fear,’ and ’sadness,’ over 50% of their emotion labels were converted to neutral in the response task.
+Our analysis reveals a systematic bias in Gemma towards rational, neutral sentiment in these response tasks.
+Notably, a portion of the original emotions still transformed into ’anger’ in the responses; for instance, ’disgust’ showed a 21% conversion rate to ’anger,’ while ’surprise’ had a 14% conversion rate.
+This suggests Gemma’s sustained sensitivity to anger-related content across both response and continuation tasks.
+Figure 3c and Figure 3d illustrate the performance of Llama 3, showing that it has the same ability to recognize and preserve emotions in continuation and response tasks, and is able to maintain the original emotional valence more consistently.
+Similarly, we see a neutral emotion bias in the Llama3 model.
+In the response and continuation tasks of the commercial model Claude, we observed that more original text emotions were more neutral in downstream tasks.
+At the same time, in the same model family, we observed that the Llama 3.3 model was also more inclined to generate neutral content in response and continuation than Llama3.
+Beyond Reddit, our analysis also included Twitter data, which revealed distinct discourse patterns surrounding climate change topic.
+Figure 4 illustrates the emotion shift between the original Twitter content and the replies generated by the LLM.
+The emotion performance patterns of the Gemma, Llama and Claude models on Twitter texts in continuation and response tasks yielded two key insights: Firstly, Gemma demonstrated heightened sensitivity to content related to anger; secondly, for models exhibited a systematic bias towards neutral sentiment when operating in interactive conversational contexts.
+third, both Claude and Llama3.3 models tend to generate content with neutral emotions.
+In addition, it is worth noting that in the continuation task of Llama3.3, its ability to continue emotions is better than Llama3.
 Figure 4: Emotional Transition Analysis of LLM Response and Continuation Tasks in Twitter Comments.
-Panels a, b, c, and d illustrate emotional transitions in content generated by Gemma and Llama models during continuation and response tasks on Twitter, respectively.
+Panels a, b, c, d e, f, g, and h illustrate emotional transitions in content generated by Gemma, Llama and Claude models during continuation and response tasks on Twitter, respectively.
 The y-axis represents the original emotions in human-authored tweets, while the x-axis shows the emotions detected in LLM-generated content.
 Each cell value represents the proportion of emotional transitions, with darker shades of red indicating higher transition frequencies.
-For example, cell values of $1.0$ would indicate complete preservation of the original emotion, while lower values indicate transitions to different emotional states.
-The heatmap reveals patterns in how each model preserves or transforms emotional content across response and continuation tasks, with notable variations in the handling of positive (e.g., joy, optimism) versus negative (e.g., anger, fear) emotions.
-In addition to Reddit, we analyzed data from Twitter, which exhibits distinct discourse patterns surrounding climate change.
-Figure 4 illustrates the emotional transitions between original Twitter content and LLM-generated responses.
-In Gemma’s continuation and response tasks, we observed that regardless of the original emotional valence—whether positive (anticipation, joy) or negative (anger, disgust)—the generated content predominantly expressed anger and anticipation.
-Among these transitions, anger represented the highest proportion, while anxiety emerged as the predominant emotion, followed by anticipation.
-A notable distinction between the two tasks was the increased frequency of responses transitioning toward anticipation.
-In Llama’s continuation task, the original emotional content similarly demonstrated a predominant shift toward anger.
-In contrast, Llama’s response task exhibited a primary shift toward anticipation.
-The performance patterns of both Gemma and Llama models across continuation and response tasks reveal two key insights: first, Gemma’s heightened sensitivity to anger-related content, and second, the models’ systematic bias toward positive affect when functioning in interactive dialogue contexts.
 Resources of LLMs’ Generated Content Emotions
 Figure 5: Emotional source analysis of LLM-generated content across platforms.
-Panels a, c, e, and g illustrate emotional transitions in Gemma and Llama models’ continuation and response tasks on Reddit data.
-Panels b, d, f, and h show corresponding results from Twitter data.
-Red lines indicate content generated from originally positive emotional text, while blue dashed lines represent content derived from negative emotional text.
-The absolute x-values represent the proportion of emotional source contributions in the generated content.
-The y-axis displays the emotional categories present in both original and generated content, while the x-axis ($-1.0$ to $1.0$) indicates the strength and direction of emotional transitions.
-We analyze the emotional sources of LLM-generated content by examining the relationship between input and output emotions.
-In Gemma’s continuation task, Figure 5a reveals that positive emotions in generated content primarily derive from positive emotional sources.
-For instance, joy-labeled content originates from anticipation (32.04%), joy (25.24%), and optimism (13.59%).
-Conversely, content expressing negative emotions predominantly stems from negative emotional sources, with anger-labeled content derived primarily from original anger (61.8%) and disgust (13.03%).
-This pattern suggests Gemma’s tendency to maintain emotional valence consistency during content continuation.
-Figure 5c illustrates Llama’s continuation task results, which differ from Gemma’s pattern.
-Most emotional content, except for optimism and joy, originates predominantly from negative emotional sources.
-While the preservation of negative emotional sources aligns with Gemma’s behavior, Llama distinctively generates positive emotional content primarily from negative emotional sources, potentially indicating an inherent positive affect bias.
-Figure 5b,d illustrate the response tasks for Gemma and Llama, respectively.
-In Gemma’s responses, most positive emotional content, except joy, originates from predominantly negative emotional sources.
-Negative emotional content maintains its source valence, with anger-labeled content derived 69.55% from original anger expressions.
-Similarly, Llama’s responses demonstrate consistent transformation of negative emotions into positive ones, exemplified by anticipation-labeled content originating 37.23% from anger.
-However, negative emotional content maintains its original valence.
-These findings demonstrate LLMs’ systematic transformation of negative emotions into positive ones during response tasks, while simultaneously exhibiting some degree of negative emotional preservation in their responses.
+Panels an illustrate emotional transitions in Gemma, Llama and Claude models’ continuation and response tasks on Reddit and Twitter data.
+The red bars indicate that the original text corresponding to the generated text is positive emotions such as joy and surprise, the blue bars indicate that the original text is negative emotions such as anger, disgust, fear and sadness, and the green bars indicate that it comes from text with neutral emotions.
+The y-axis displays the emotional categories present in both original and generated content.
+We analyzed the emotional sources of LLM-generated content by examining the relationship between input and output emotions.
+As shown in Figure 5a, the results show that Gemma did not simply copy or continue the original emotion of the input text, but perform significant emotion modulation.
+A core finding is that the model has a general tendency to moderate negative initial emotion.
+Specifically, when the original text carries negative emotion (as shown by the blue bar in the figure 5, representing anger, disgust, fear, or sadness), a considerable proportion of the generated content emotion will turn into neutral (source shown by the green bar) or even positive emotion (source shown by the red bar, representing joy or surprise).
+For example, in the continuation and response tasks of multiple models (such as Gemma, Llama, Claude), as well as on the Reddit and Twitter datasets, it can be observed that the generated text with ”neutral” or ”joy” emotion has a significant negative emotion contribution in its original emotion input.
+In addition, the model also shows a general trend towards neutral emotion, that is, the generated neutral emotion text comes from a wide range of original positive, negative, and neutral inputs.
+Even if the original input is positive, the generated content has a certain probability of being adjusted to neutral.
+These emotion conversion patterns show high consistency between different LLMs, task types (text continuation and response), and data sources (Reddit and Twitter), suggesting that this may be an inherent characteristic or common training result of current mainstream LLMs in emotion processing.
 Comparative Analysis of Emotional Intensity between LLMs and Human Text
-Beyond examining emotional transitions, we investigated the quantitative differences in emotional intensity between LLM-generated and human-authored content.
-This analysis specifically focused on determining whether LLM-generated content exhibits higher or lower emotional intensity compared to human expressions.
-To quantify emotional content, we employed a probabilistic model that assigns normalized scores (0 to 1) to each emotional category within the text.
-These probability values were interpreted as emotional intensity scores (Miyazaki et al.
+We analyzed the differences in emotional intensity between LLM-generated and human-authored content, focusing on whether LLMs exhibit higher or lower emotional intensity.
+Emotional content was quantified using a probabilistic model assigning normalized scores (0 to 1) to each emotional category, interpreted as intensity scores (Miyazaki et al.
 2024).
-We categorized emotional expressions into five distinct intensity groups based on these scores.
-Statistical analysis comprised an ANOVA test to evaluate differences in emotional intensity across groups, followed by Tukey’s post-hoc test to identify significant pairwise variations(Elnaggar, Mohamed, and Gehan 2024).
-Table 1: Test for significant differences in affective values between groups
+These scores were grouped into five intensity levels.
+Statistical analysis included ANOVA to compare group differences and Tukey’s post-hoc test for significant pairwise variations(Elnaggar, Mohamed, and Gehan 2024).
+Table 1: ANOVA test of different model emotions
 Platform Emotions F statistic P value
-Reddit anger 197.316 6.55E-166
-anticipation 80.899 1.18E-67
-disgust 35.307 4.12E-29
-optimism 59.526 6.83E-50
-fear 45.168 2.70E-37
-sadness 18.736 3.58E-15
-Twitter anger 384.285 9.5078E-320
-anticipation 16092.479 0
-disgust 179.181 1.80E-143
-optimism 6.088 6.95E-05
-fear 73.936 6.99E-60
-joy 1742.163 0.00E+00
-sadness 148.411 6.85E-98
-Analysis of variance (ANOVA) results presented in Table 11 indicate statistically significant differences ($P<0.05$) across the five groups in emotional intensity values for anger, anticipation, disgust, optimism, fear, joy, and sadness on Twitter.
-Similar significant variations were observed in the Reddit dataset for anger, anticipation, disgust, optimism, fear, and sadness.
+Reddit anger 114.2381 <0.001
+joy 80.899 <0.001
+disgust 29.8564 <0.001
+surprise 24.9967 <0.001
+fear 17.3482 <0.001
+sadness 5.5149 <0.001
+neutral 325.2502 <0.001
+Twitter anger 202.1090 <0.001
+joy 34.4011 <0.001
+disgust 22.4241 <0.001
+surprise 51.5744 <0.001
+fear 75.0905 <0.001
+sadness 9.2834 <0.001
+neutral 552.6669 <0.001
+Analysis of variance (ANOVA) results presented in Table 1 indicate statistically significant differences ($P<0.01$) across the nine groups in emotional intensity values for anger, disgust, fear, sadness, joy, surprise and neutral on Twitter.
+Similar significant variations were also observed in the Reddit dataset for the seven emotions mentioned above.
 Tukey’s post-hoc analysis identified several significant differences across three comparison categories: within-model, between-model, and model-to-human comparisons.
-On Twitter discussions of climate change, within-model analyses revealed that Gemma’s continuation content exhibited significantly lower anticipation values compared to its response content.
-Similarly, Llama’s continuation content showed significantly lower anticipation values than its response content.
-Between-model comparisons demonstrated that Gemma’s continuation content expressed significantly higher anticipation values than Llama’s continuation content.
-Additionally, Gemma’s response content showed significantly higher anticipation values compared to Llama’s responses.
-In model-to-human comparisons, both models’ generated content (continuation and response) demonstrated significantly lower anticipation values compared to the original human text.
-These findings demonstrate that: (1) both models express higher anticipation values in response tasks compared to continuation tasks; (2) Gemma consistently generates content with higher anticipation values compared to Llama across both tasks; and (3) both models generate content with significantly reduced anticipation values compared to human-authored text, indicating a systematic reduction in emotional intensity during the generation process.
-Table 2: Tukey’s post-hoc test of Reddit emotions
-Emotion Within Group1 Between Groups2 Comparison with original
-Gemma Con
-vs. Resp
-Llama Con
-vs. Resp
-Gemma Con
-vs. Llama Con
-Gemma Resp
-vs. Llama Resp
-Gemma Con
-vs Original
-Gemma Resp
-vs Original
-Llama Con
-vs. Original
-Llama Resp
-vs. Original
-Anticipate <*** <*** <*** <***
-Joy <**
-Disgust <*** <** >*** <*** >* <***
-Sadness <*** >* <*** <*** <***
-Anger <*** <*** <*** <*** <*** >***
-Optimism >*** >** <*** <*** <*** <***
-Fear <*** <*** <*** <*** <***
+Each emotion score represents its degree of emotion, and the higher the score, the stronger the emotion.
+As shown in the table2, intra-model comparisons indicate that emotional scores for certain emotions—such as anger and fear—are generally higher in continuation tasks than in response tasks.
+This suggests that LLMs tend to adopt a more neutral emotional tone when operating as conversational agents.
+Furthermore, when compared to the original human-authored texts,shown in the table3, most LLM-generated outputs exhibit lower emotional intensity, particularly in dimensions such as anger, fear, disgust, and joy.
+This may indicate that LLMs express emotions in a more restrained and subdued manner relative to humans.
+Cross-model comparisons also reveal that emotional expression varies significantly depending on task type and data source.
+For instance, within the Llama model family, Llama 3 demonstrates higher emotional scores than Llama 3.3, suggesting a greater tendency toward heightened emotional expression.
+Table 2: Tukey’s post-hoc test of LLM-generated content emotion values
+Emotion Within Group1 Between Groups2 Platform
+Gemma Con vs Resp Llama Con vs Resp Claude Con vs Resp Llama3.3 Con vs Resp Gemma Con vs Llama Con Gemma Con vs Claude Con Llama Con vs Claude Con Llama Con vs Llama3.3 Con Gemma Resp vs Llama Resp Gemma Resp vs Claude Resp Llama Resp vs Claude Resp Llama Resp vs Llama3.3 Resp
+anger >*** <*** >*** - >** - >*** - <*** <* <*** >*** Reddit
+fear <*** - <** <*** <*** >*** - >* - - - -
+sadness >** <*** <** - >*** - >*** - <** >*** >*** -
+joy - - >*** >*** >*** - >*** - >* <** - >***
+surprise - >** - - >*** <** - - >*** <* - -
+disgust <* <*** <** <*** - - - >*** - - - -
+neutral >*** >*** >** >*** - <*** <*** >*** <*** >* <*** >***
+anger - - >*** >*** >*** - <*** - >* >** - >*** Twitter
+fear >*** >*** >** >*** - >*** >*** >*** <*** <*** >*** >***
+sadness <*** - <* <*** <*** <*** - >* - - - -
+joy - >*** - - >*** >** - - >*** >* - -
+surprise >* <*** <** - >*** - <*** - <** <*** <*** -
+disgust >*** <*** >*** - >** - <*** - <*** >* >*** >***
+neutral <* <*** <** <*** - - - >*** - - - -
   • 1
@@
   • 4
-    The symbolic $<$ and $>$ indicate that the data in the previous column is less than or greater than the data in the next column.
-The “-” means there is no statistically significant difference between the two columns of data.
-Table 3: Tukey’s post-hoc test of Twitter emotions
-Emotion Within Group1 Between Groups1 Comparison with original1
-Gemma Con
-vs. Resp
-Llama Con
-vs. Resp
-Gemma Con
-vs. Llama Con
-Gemma Resp
-vs. Llama Resp
-Gemma Con
-vs Original
-Gemma Resp
-vs Original
-Llama Con
-vs. Original
-Llama Resp
-vs. Original
-Anticipate <*** <*** >*** <*** <*** <*** <*** <***
-Joy <*** <*** <*** <***
-Disgust <*** <*** <* <*** <***
-Sadness <*** <*** <*** <*** <*** <***
-Anger <*** <*** <*** <*** <*** <*** <*** <***
-Optimism <*** <*** <*** <***
-Fear <*** <*** <** <** <*** <*** <*** <***
+    The symbolic $<$ and $>$ indicate that the data in the previous column is less than or greater than the data in the next column. The “-” means there is no statistically
+    significant difference between the two columns of data.
+Table 3: Tukey’s post-hoc test of LLM-generated content and human text emotion values
+Platform Emotions Gemma Con vs Original1 Gemma Resp vs Original Llama Con vs.
+Original Llama Resp vs.
+Original Claude Con vs Original Claude Resp vs Original Llama3.3 Con vs Original Llama3.3 Resp vs Original
+Reddit anger <*** <*** <*** <*** <*** <*** <*** <***
+fear <*** - - - - - <*** -
+sadness - - - - >* >*** - -
+joy - - <*** - - <** <** <***
+surprise - - <*** <*** - - <*** -
+disgust >*** >*** >*** >*** >** >*** - >***
+neutral >*** <*** >*** <** <*** <*** - <***
+Twitter anger - - <*** - - <** <** <***
+fear >*** <*** >*** <** <*** <*** - <***
+sadness <*** - - - - - <*** -
+joy - - <*** <*** - - <*** -
+surprise - - - - >* >*** - -
+disgust <*** <*** <*** <*** <*** <*** <*** <***
+neutral >*** >*** >*** >*** >** >*** - >***
   • 1
-    The comparison of the emotion values of continuation and response content generated by the same model.
+    The comparison of the emotion values of the LLM-generated content by different model and human text.
   • 2
-    The comparison of the emotion values of the same task but generated by different model.
-  • 3
-    The stars indicate the p values of the Mann–Whitney U test: *** for $p<0.001$, ** for $p<0.01$, and * for $p<0.05$.
+    The stars indicate the p values of the Tukey’s post-hoc test: *** for $p<0.001$, ** for $p<0.01$, and * for $p<0.05$.
   • 4
-    The symbolic “$<$ and $>$” indicate that the data in the previous list is less than or greater than the data in the next list.
-The “-” means there is no statistically significant difference between the two columns of data.
-Further analysis of additional emotional dimensions revealed that Tukey’s post hoc test results indicate significant differences ($P<0.05$) between LLM-generated and original texts across multiple emotional categories.
-Both Gemma and Llama’s generated content, in both continuation and response tasks, exhibited significantly lower intensities of positive emotions, particularly joy and optimism.
-A nuanced pattern emerged in the expression of negative emotions.
-The continuation texts generated by both models demonstrated attenuated levels of sadness, anger, disgust, and fear compared to their respective response texts and the original human content.
-Notably, Gemma’s response text exhibited significantly lower intensities of anger, disgust, and fear compared to Llama’s responses.
+    The symbolic “$<$ and $>$” indicate that the data in the previous list is less than or greater than the data in the next list. The “-” means there is no statistically
+    significant difference between the two columns of data.
 These findings suggest two key insights: first, LLMs demonstrate systematic suppression of certain negative emotions, particularly in continuation tasks; second, the response task appears to operate under distinct generative mechanisms, resulting in differential emotional expression patterns.
 Furthermore, the consistent reduction in optimism across all LLM-generated texts relative to human-authored content indicates a systematic constraint in LLMs’ capability to fully capture and convey positive emotional states.
-Semantic Coherence Analysis between LLM-Generated and Human-Authored Content
-In LLM-human interactions, beyond examining emotional congruence and intensity patterns demonstrated in our previous experiments, we investigated the models’ capacity to maintain topical coherence and generate contextually relevant responses.
-To quantify this relationship, we employed cosine similarity as a metric to assess semantic alignment between generated and original content.
-This approach provides a quantitative framework for evaluating the semantic fidelity of LLM-generated responses across different interaction contexts.
+Evaluating Semantic Consistency of LLM-Generated Content in Social Media Contexts
+In LLM-human interactions, we analyzed models’ ability to maintain topical coherence and contextual relevance using scores labeled by LLM-as-judge (Zheng et al.
+2023) to measure semantic and logic alignment with the original content, providing a framework for evaluating semantic and logic fidelity across contexts.
 a
 b
-Figure 6: Distribution of cosine similarity scores for LLM-generated content, visualized using violin plots comparing Gemma and Llama models’ continuations and responses across Reddit (panel a) and Twitter (panel b) platforms.
-Cosine similarity values range from -0.2 to 1.0, with median scores indicated numerically for each distribution.
-Table 4: Mann-Whitney U test for cosine similarity between LLM-generated content and two platform text
-Comparison
-group
-Statistics P value
-Reddit Inner
-Gemma
-continuation
-vs
-Gemma
-response
-56740052 0.999999995
-Llama
-continuation
-vs
-Llama
-response
-66851679 4.05E-58
-Inter
-Gemma
-continuation vs
-Llama
-continuation
-68723025 8.90041E-90
-Gemma
-response vs
-Llama
-response
-78292192 0.00E+00
-Twitter Inner
-Gemma
-continuation vs
-Gemma
-response
-93014920 9.74E-251
-Llama
-continuation vs
-Llama
-response
-60081284 1
-Inter
-Gemma
-continuation vs
-Llama
-continuation
-118641175 0
-Gemma
-response vs
-Llama
-response
-69518835 1
-  • 1
-    The $p<0.05$ indicate that the data in the previous group is significantly higher than the data in the latter group.
-Table 4 presents the Mann-Whitney U test results across the four experimental conditions.
-The analysis reveals that both Gemma and Llama maintained substantial semantic alignment with the original text in their continuation and response tasks.
-Figure 6a illustrates that semantic similarity values on the Twitter platform predominantly exceeded 0.5, indicating strong semantic coherence between LLM-generated content and input text.
-This suggests the models’ capability to comprehend and generate contextually relevant responses while maintaining topical coherence.
-Similar patterns emerged in the Reddit dataset, as shown in Figure 6b, where all four sets of cosine similarities demonstrated values above 0.5.
-The Mann-Whitney U test (Table 4) revealed significant within-model and cross-task variations in semantic similarity.
-On the Twitter platform, Gemma’s continuation task demonstrated significantly higher semantic similarity compared to its response task, while Llama exhibited the opposite pattern, with responses showing significantly higher similarity ($P<0.01$).
-Cross-model analysis revealed task-specific differences: Gemma’s continuation task demonstrated significantly higher semantic similarity compared to Llama’s continuation task ($P<0.01$), while Llama achieved higher similarity scores in the response task.
-On Reddit, Llama’s continuation content exhibited significantly higher semantic similarity compared to its response content ($P<0.01$).
-Furthermore, Gemma’s generated content demonstrated significantly higher semantic similarity compared to Llama’s in both continuation and response tasks ($P<0.01$).
+Figure 6: Distribution of faithfulness and consistency scores for LLM-generated content, visualized using stacked histograms comparing Gemma, Llama, and Claude models’ continuations and responses across Reddit (panel a) and Twitter (panel b) platforms.
+The x-axis denotes rating categories, with 5 representing ’Excellent’, 4 ’Good’, 3 ’Fair’, 2 ’Poor’, and 1 ’Very Poor’.
+Figure 6 plotted the frequency of each score (ranging from 1 to 5) separately for each task.
+This allows us to examine whether certain tasks tend to produce more consistent model outputs.
+The results suggest that while most tasks concentrate around higher scores, there are subtle differences in how each subtask is rated by the evaluator.
+We find that most models are capable of generating high-quality content in both response and continuation tasks, indicating that current LLMs possess strong capabilities in understanding and interacte with human language.
+However, subtle differences remain across models.
+For instance, the Llama 3.3 model demonstrates consistently strong performance in content generation across both types of datasets.
+Comparized with Llama 3’ result, this may also indicate to some extent the impact of different model sizes on the quality of generated content.
+Additionally, model performance varies across platforms.
+For example, the Claude model shows higher fidelity and consistency in its outputs on the Reddit dataset, which may be attributed to the more complex linguistic environment of Twitter.
 Discussion
-This study investigated the consistency between LLM-generated content and human input, with particular focus on emotional content transmission.
-Given that text often encodes substantial emotional information, we hypothesized that LLMs may demonstrate diverse emotional patterns.
-As LLMs increasingly integrate into human social contexts through chatbots and social agents, understanding their semantic consistency and emotional transmission patterns becomes crucial for human-machine interaction.
-We decomposed the fundamental question of LLM emotional capability into three specific research questions, based on our empirical findings.
-Several key insights emerged from our analysis.
-In the continuation task, Gemma demonstrated a systematic transformation of most original emotions toward anger, indicating a bias toward negative emotional expression, particularly in anger intensification.
-However, Gemma exhibited capacity for emotional preservation, specifically for optimism and surprise, suggesting ability to perceive and maintain certain emotional valences through the continuation process.
-Conversely, the Llama model demonstrated enhanced capability in preserving original emotional states during continuation tasks, particularly for anger, anticipation, fear, optimism, and sadness, with relatively lower emotional transformation rates, indicating superior emotional continuity (RQ1).
-We posit that during continuation tasks, LLMs employ emotion-specific lexical markers, syntactic patterns, and contextual cues to maintain emotional consistency with source texts.
-Emotions characterized by prominent semantic features (such as anger, anticipation, and sadness) typically manifest through distinct emotional vocabulary and strong affective markers, facilitating model recognition and continuation.
-Moreover, the semantic salience of anger may enhance its recognition and preservation in generative tasks, while positive emotions like optimism or joy may require more robust contextual support, making them susceptible to disruption by LLM-generated negative content.
-In the response task, both Gemma and Llama models exhibited a systematic bias toward positive emotions, predominantly converting original emotions into anticipation and optimism.
-This pattern suggests an embedded bias from the training process favoring positive affect.
-However, the persistent presence of anger in a subset of responses demonstrates these models’ susceptibility to negative emotional content during the response generation process.
-Our experimental findings indicate that both Gemma and Llama models exhibit emotional capabilities and can effectively recognize and maintain human emotional states in continuation tasks.
-This emotional recognition capability is fundamental for LLMs, as it underlies their ability to comprehend both user requirements and emotional context.
-Regarding emotional intensity (RQ2), our results demonstrate that LLM-generated content exhibits significantly attenuated intensity across most emotional dimensions compared to human-authored texts.
-This suggests that LLMs exhibit more moderated emotional expression, maintaining more constrained emotional ranges across both positive and negative affects.
-Furthermore, the consistently lower emotional intensity values in both continuation and response tasks, compared to original texts, indicates potential limitations in LLMs’ capacity to fully capture and convey emotional depth.
-Analysis of emotional types reveals distinct patterns in LLM-generated content.
-Regarding positive emotions, LLM-generated texts demonstrate significantly lower intensities of joy and optimism compared to human-authored texts, suggesting limitations in the models’ capacity to fully comprehend and express positive emotional states characteristic of human experience.
-Similarly, for negative emotions, LLM-generated content exhibits attenuated intensities of sadness, anger, disgust, and fear, indicating systematic suppression of negative emotional expression in downstream tasks.
-In the context of global challenges such as climate change, which transcends temporal, spatial, and geographical boundaries and impacts long-term human development, public discourse exhibits diverse emotional responses.
-As social media platforms rapidly evolve, they become crucial channels for public communication during climate-related events, such as the Australian bushfires.
-The effective management of public emotions during critical periods following such events represents a key component of public opinion governance.
+This study utilized the Twitter and Reddit datasets on climate change to systematically evaluate three aspects of LLM-generated text versus original social media posts: (1) emotional consistency in both continuation and response tasks (and differences in different models), (2) emotional intensity across those tasks, and (3) logic similarity and context-awareness between model outputs and original social media texts.
+By combining these analyses, we aim to uncover the potential effects on emotional dynamics that LLMs have when implemented in online discussion.
+In the continuation task, Gemma consistently transformed most emotions toward anger, showing a bias toward negative expression and intensifying anger.
+However, it maintained optimism and surprise, suggesting some ability to preserve emotional valence.
+On the other hand, Llama better preserved original emotions like anger, anticipation, fear, optimism, and sadness, with fewer emotional shifts, reflecting stronger emotional continuity (RQ1).
+By combining the comparative analyses of emotional consistency and intensity, this study demonstrates that in both response and continuation tasks, LLMs systematically shift emotional content toward neutrality and lower its intensity across most emotional categories compared to original texts.
+Particularly, generated outputs display a higher prevalence and intensity of “neutral” manner, especially in response tasks rather than continuation tasks.
+This pattern shows that LLMs do have effect on the change of emotion when engaged in online discussion, and they’re turning the discussion in a emotionally-moderate way.
+As aforementioned, emotion plays a critical role in navigating social media discussion, especially in controversial topics like climate change.
+While the climate change discussion online is quite polarized (Treen et al.
+2022b), such a tendency to neutrality could help defuse heated exchanges and foster calmer, less contentious dialogue.
+Hence, even though LLMs cannot keep the emotional consistency as human, its preferences toward moderation could benefit the highly polarized and emotionally debatable topics online, and lead the discussion more neutrally.
+However, beyond this “neutral” tendency, what should be aware is that the negative emotions like “anger” and “fear” also showed increase in LLM-generated continuations and responses in most of the cases, which indicates that these models can amplify such emotional cues.
+Anger and fear are among the most salient emotional frames in climate discourse (Nabi, Gustafson, and Jensen 2018; Davidson and Kecinski 2022), and are significantly associated with climate activism, engagement, and has great effect on people’s attitudes and behaviors (Stanley et al.
+2021; Gregersen, Andersen, and Tvinnereim 2023).
+An uptick in fear and anger in LLM outputs may therefore shape users’ perceptions of climate change.
+Considering the echo chamber effect on social media, LLMs’ propensity of fear and anger could increase people’s perceived threat by climate change, which in turn leads to higher awareness and action intentions towards climate crisis.
+In terms of semantic consistency, the results of LLM-as-judge (Figure 6) show that LLMs generated text could generally keep the consistency to social media texts.
+This comparative findings proves a critical evidence from the perspective of emotion that LLMs could generally “understand” the context of a social media post and produce continuations or replies that align both thematically and logically with the source content.
+However, the consistency varies due to different types of LLMs, while most of the scores are high, some specific model like Claude’s performances are dispersed, featuring a nontrivial tail into other scores.
+It indicates that some models may be more sensitive to prompt phrasing or exhibit less robust world-knowledge integration.
+The findings from semantic consistency imply that LLMs nowadays can generally preserve semantic and logic coherence in social-media dialogue, which makes the LLMs-generated content hard to distinguish on social media.
+Taken together, the emotional and semantic consistency examined in this study could be seen as a technological improvement of context-awareness in LLMs’ text generation capabilities.
+However, it could also raise practical concerns to general social media users for the risks of LLMs being manipulated for some malicious or intentional purposes, such as misinformation.
+Effectively managing public emotions during controversial discussion online is crucial for governance of public opinion.
 Recent advances in Artificial Intelligence Generated Content (AIGC), driven by Generative AI (GAI) technology, have garnered attention beyond computer science (Cao et al.
@@
 Given the increasing integration of LLMs into daily life, their emotional characteristics significantly influence opinion leadership, as emotional content shapes public perception and discourse framing.
-Our experimental results (Figure 3 and Figure 4) demonstrate that LLMs systematically transition toward positive emotions during response tasks.
-This positive bias suggests potential utility in public opinion events, offering constructive responses and mitigating negative public sentiment.
-However, the observed patterns of negative emotional transmission pose potential societal risks.
-LLMs’ capability to generate contextually relevant content could be exploited for malicious purposes.
-Beyond concerns regarding misinformation dissemination, the potential intensification of emotional polarization presents a significant challenge in LLM deployment.
-From the perspective of human-computer interaction, emotional support serves as a fundamental component in enhancing social interactions, facilitating psychological interventions, and improving customer service outcomes through addressing emotional needs.
-The quality of emotional support and user understanding significantly impacts long-term user engagement and trust in LLM interactions (Schneider, Flores, and Kranz 2024).
-Addressing our third research question, we analyzed embedding representations of generated and original content, employing cosine similarity as a metric for semantic proximity.
-Our findings indicate that both Gemma and Llama consistently demonstrate high cosine similarity values, suggesting robust capability in capturing and reproducing semantic features.
-Models have shown a remarkable ability to represent, comprehend, and generate human-like text.
-Compared to prior Natural Language Processing (NLP) approaches, one of the most striking advances of LLMs is their ability to generalize their “knowledge” to novel scenarios, contexts, and tasks(Peters and Matz 2024) Future implications suggest that while LLMs demonstrate competence in generating semantically coherent content, opportunities exist for enhancing alignment between generated outputs and nuanced human context.
-Future research directions should explore mechanisms for refining LLMs’ comprehension of implicit meaning and contextual subtleties, thereby enhancing user experience and expanding application domains.
+Future implications suggest that while LLMs generate semantically coherent content, there is potential to improve alignment with nuanced human contexts.
+Research should focus on refining LLMs’ understanding of implicit meaning and contextual subtleties to enhance user experience and broaden application domains.
 Limitation and Future Work
-This study contributes to understanding the emotional dynamics of human-AI interactions, while suggesting several avenues for future research based on current limitations.
-First, regarding experimental design, our study was limited to data from Reddit and Twitter platforms, potentially under-representing the broader social media ecosystem.
-Other platforms such as YouTube, Instagram, and TikTok exhibit distinct user behaviors and content structures (Hilde A.
+This study enhances understanding of emotional dynamics in human-AI interactions while highlighting opportunities for future research to address current limitations.
+First, regarding experimental design, our analysis relied on Reddit and Twitter data.
+Platforms like YouTube, Instagram, and TikTok exhibit unique user behaviors and content structures (Hilde A.
 M.
 Voorveld and Bronner 2018).
-Furthermore, our analysis was bounded by the capabilities of open-source emotion models (Llama and Gemma), which may demonstrate different performance characteristics compared to proprietary commercial models like ChatGPT or Google’s Bard.
-Future research could expand the scope by incorporating data from additional social media platforms and evaluating commercial models to enhance generalizability.
-Second, our experimental scope, focused on English-language climate change discourse, provides limited insight into model performance across diverse topics and languages.
-The emotional expression patterns of LLMs may vary significantly across different subject domains.
-Future research directions could incorporate multilingual datasets to identify cross-linguistic semantic variations (Zhao et al.
-2020) and investigate potential biases within LLMs across isolated or intersecting semantic spaces (Hassan et al.
-2018).
-Third, while we utilized social media to obtain human-authored content, the dataset potentially includes automatically generated content and social bot interactions that may not accurately represent genuine human expression.
-Future research should implement more rigorous data validation protocols to enhance dataset quality and control for confounding variables.
+Second, the explaination for the findings in emotional inconsistency of LLMs requires further exploration.
+While we revealed that emotions vary both across different LLMs and across generation tasks, for example, the “neutral” emotion take the major role among all emotions in all cases but differs between continuation and response tasks and LLMs, likewise, the overall proportion of “anger” significantly increases in all cases, and different models exhibit distinct emotional profiles.
+However, because the training data and internal weights of these models are not publicly available, we cannot at present pinpoint the underlying causes of these differences—whether they stem from pretraining data biases, fine-tuning strategies, or discussion context.
+Future work should work on the interpretability analyses, controlled comparisons, or replication on open datasets, to investigate the mechanisms driving these emotion changes.
+Third, this study may have potential data contamination due to the use of publicly available data from Twitter and Reddit.
+Twitter and Reddit data are publicly accessible and may overlap with the LLMs’ pre-training corpora, model outputs could partly reflect memorized patterns rather than genuine reasoning in real-world contexts, which limits the generalization of the results.
+While we cannot fully quantify the contamination, we acknowledge its impact and encourage future research to explore methods for mitigation—such as by using post-training data that is demonstrably out-of-scope of LLM training sets, or by focusing on data published after known pretraining cutoff dates.
+Besides, social media-derived content may include automated posts and social bot interactions, potentially skewing results.
+Future research should adopt stricter data validation protocols to ensure datasets quality and reduce confounding variables.
+Fourth, since human expert evaluation is expensive, we adopt a hybrid evaluation strategy that combines human judgment with LLM assistance based on a small but diverse sample across platforms, tasks, and models.
+While this approach achieves high-quality and interpretable scores, we acknowledge that the limited sample size may not capture the full distribution of consensus behaviors across the datasets.
+Future work could extend this approach, such as scalable pure LLM evaluation and automatic divergence detection for selective human review.
 Conclusion
-This study examined the semantic and emotional consistency of content generated by LLMs.
-Recognizing the emotional information embedded in text and the way LLMs handle these emotions is essential as they become more involved in human social contexts.
-Our findings revealed that Gemma demonstrated a tendency to amplify negative emotions, particularly anger.
-Despite this bias, Gemma effectively perceived and transferred the original emotional tone of the text.
-On the other hand, the Llama model displayed stronger emotional retention across a broader spectrum, including anger, expectation, fear, optimism, and sadness, with fewer transitions to other emotions.
-For the semantic problem, both models show better performance for both continuation and response.
-LLMs can understand human input to some extent.
-These observations highlight distinct emotional dynamics between LLMs, offering insights for enhancing their design and application in emotion-sensitive contexts.
-Acknowledgments
-This study is partially supported by the 2023-2024 Annual Seed Research Funding of the Department of Science and Technology of Communication, University of Science and Technology of China (#ASRF2023).
+In this study, we evaluated how large language models handle emotional and semantic content in social-media contexts.
+By comparing four representative models—Gemma, Llama3, Claude, Llama3.3—on both continuation and response tasks, we uncovered distinct patterns in their emotional behaviors and contextual understanding.
+Across both continuation and response tasks, LLMs tended to shift emotional content toward neutrality and generally reduced emotional intensity.
+This effect was more pronounced in response tasks, suggesting that LLMs favor emotionally moderate language when addressing existing posts.
+LLM-as-judge evaluations revealed that generated texts largely preserved thematic and logical coherence with source texts.
+While most models scored consistently high, variability in some (e.g., Claude) highlights sensitivity to prompt phrasing and model-specific robustness.
+The neutrality bias of LLMs could help defuse polarization and foster calmer dialogue around contentious topics.
+Conversely, the amplification of anger and fear poses both opportunities such as mobilization and risks like manipulation or echo-chamber reinforcement.
+These findings reveal key emotional dynamics between LLMs, offering insights for improving their design and use in emotion-sensitive applications.